OpenStudy (anonymous):

3x + 2y = 6 4x + y = 1 Solve the system of equations

OpenStudy (anonymous):

from equation 2: y = 1 - 4x

OpenStudy (anonymous):

substitute y= 1- 4x in equation 1: 3x + 2(1-4x) = 6

OpenStudy (anonymous):

simplifying will give us 3x + 2 -8x = 6 -5x = 4 x = -4/5 = -0.8

OpenStudy (anonymous):

now substitute x = -4/5 either to equation 1 or 2: then you will get the value of y

OpenStudy (zzr0ck3r):

There is another way, and I think it is often easier \(3x+2y=6\\4x+y=1\) Multiply the second equation by \(-2\) and we get \(3x+2y=6\\-8x+-2y=-2\). Now add the two equations together. \(-5x=4\implies x=-\frac{4}{5}\) Now plug this into anything with an \(x\) and \(y\) to solve for \(y\).